OpenStudy (anonymous):

the side of a cube is expanding at a constant rate of 6 inches per second. What is the rate of change of the volume, in inches^3 per second, when the total surface area of the cube is 54 inches^2?? Help.!

OpenStudy (anonymous):

ds/dt = 6 in/sec S.A = 6s^2 54 = 6 s^2 54/6 = s^2 9 = s^2 s = 3 volume = s^3 dv/dt = 3s^2 ds/dt dv/dt = 3*3^2*6 dv/dt = 162

OpenStudy (anonymous):

what does the s stand for in ds/dt. i know that the t stands for time

OpenStudy (anonymous):

change in side with respect to time let s = side

OpenStudy (anonymous):

oh okay i got you now.

OpenStudy (anonymous):

can you verify the answer ?